OpenStudy (anonymous):

How can you find the value of 'a' in a quadratic function (y=ax+bx+c) when given one co-ordinate of the x intercept and the y intercept?

hartnn (hartnn):

what are the intercepts and point given ?

hartnn (hartnn):

ok, let me give you an example : if x -intercept is 2,0 and y-intercept is 0,4 so, we have f(x) = ax^2 + bx +c to get y-intercept you put x = 0 so, y-intercept = f(0) = 0+0+c = c so, c= 4 to get x intercept, put f(x) =0 ax^2+bx +4 = 0 also, since we know that 2,0 lies on f(x), we can put x=2 4a+2b +4= 0 you need some more info to get 'a' are both x-intercepts given ? or only one ?